# An End-to-End Room Geometry Constrained Depth Estimation Framework for Indoor Panorama Images

### v1

@@ Abstract
Predicting spherical pixel depth from monocular $360^{\circ}$ indoor panoramas is critical for many vision applications. However, existing methods focus on pixel-level accuracy, causing oversmoothed room corners and noise sensitivity. In this paper, we propose a depth estimation framework based on room geometry constraints, which extracts room geometry information through layout prediction and integrates those information into the depth estimation process through background segmentation mechanism. At the model level, our framework comprises a shared feature encoder followed by task-specific decoders for layout estimation, depth estimation, and background segmentation. The shared encoder extracts multi-scale features, which are subsequently processed by individual decoders to generate initial predictions: a depth map, a room layout map, and a background segmentation map. Furthermore, our framework incorporates two strategies: a room geometry-based background depth resolving strategy and a background-segmentation-guided fusion mechanism. The proposed room-geometry-based background depth resolving strategy leverages the room layout and the depth decoder’s output to generate the corresponding background depth map. Then, a background-segmentation-guided fusion strategy derives fusion weights for the background and coarse depth maps from the segmentation decoder’s predictions. Extensive experimental results on the Stanford2D3D, Matterport3D and Structured3D datasets show that our proposed methods can achieve significantly superior performance than current open-source methods. Our code is available at https://github.com/emiyaning/RGCNet.
@@ I Introduction
With the advent of consumer-grade omnidirectional cameras such as the Ricoh Theta, Samsung Gear 360, and Insta360, the acquisition of panoramic images has been significantly simplified. The $180^{\circ}\times 360^{\circ}$ field of view offered by panoramas renders them particularly valuable for indoor 3D perception applications[1]. To enable 3D understanding of indoor environments from a single omnidirectional image, depth estimation becomes a fundamental requirement[2, 3, 4, 5, 6, 7, 8]. This potential for inferring entire scene structure from a single panorama has motivated active research on panoramic depth estimation.
Figure 1: 3D visualization of panorama depth estimator’s predictions on the Stanford2d3d dataset. The left side shows the ground-truth visualization, middle column shows the visualization of Panoformer’s prediction, right side shows the visualization of our framework’s prediction.
However, the ultra-wide field of view in panoramic imaging inherently introduces object distortion. To address this issue, current panoramic depth estimation approaches primarily employ two kinds of pathway[9]. One strategy designs dedicated feature extractors motivated by panoramic imaging principles to handle distortion directly[10, 11]. The other strategy projects the equirectangular panorama (ERP) into six cubemap faces, employs feature encoders separately on both the ERP and cubemap projections [12], fuses their features, and subsequently predicts depth. These existing methods have achieved remarkable achievements in the field of depth estimation of panoramic images. However, current panoramic depth estimation methods excessively prioritize local pixel accuracy over geometric room structures, leading to two issues: (1) Noisy ground truth data induces local overfitting that propagates errors to adjacent regions; (2) Local correlations cause inaccurate predictions at 3D discontinuity regions (e.g., wall corners). As described in Fig. 1, the pre-trained panorama depth estimator PanoFormer[10] shows limited performance on the wall corners while easily be disturbed by noisy ground truth data. These two problems caused by ignoring room geometry seriously hinder the application of existing depth estimation models to real-world indoor scenarios.
To address existing problems in panoramic depth estimation, this paper proposes a depth estimation framework named as room geometry constrained depth estimation network (RGCNet). The proposed RGCNet based on multi-task learning unifying three complementary tasks: depth estimation, room layout estimation, and background segmentation. This architecture employs a shared panoramic encoder to extract multi-scale features, with three dedicated decoders generating: coarse depth map, room layout map, and background segmentation mask from these features. The depth estimation branch retains the architecture of existing high-performance depth estimators [10, 11]. However, our RGCNet treats the output of the depth decoder as a first-stage prediction and refines it using the room geometry constrain which generated from room layout and background segmentation. In our framework, the initial depth estimate and room layout are jointly utilized to compute the background depth maps. By incorporating geometric constraints of the room, these derived background depth maps are not only more robust to noise in the ground-truth depth but also yield higher accuracy in structured regions such as walls and corners. Furthermore, RGCNet leverages an end-to-end multi-task framework to decode the background depth and fuse it with the coarse depth prediction in a single efficient step. To generate the background map, a proposed room-geometry-based background depth resolving strategy leverages the layout and depth decoder’s output to calculating the corresponding background depth. To fuse the background and depth decoder’s prediction, a background-segmentation-guided fusion strategy derives fusion weights for the background and coarse depth maps from the segmentation decoder’s predictions. In general, our contributions can be summarized into the following three points:
  • A room-layout constrained depth estimation framework RGCNet has been proposed to use room structural geometry to get corrected depth predictions.
  • A background depth map resolved strategy has been proposed, which extract the room structural geometry information from multi-task decoder’s prediction.
  • A adaptively fusion strategy has been proposed, which adaptively based on extracted the geometry information to refine the depth prediction.
@@ II-A Single Projection Inputs
Panoramic images employ spherical representations with $180^{\circ}$ vertical and $360^{\circ}$ horizontal fields of view. Processing typically involves projecting these spherical images onto 2D planes through perspective mapping. Common modalities include equirectangular[13, 14, 10, 11, 15, 16, 17], tangent[18, 19, 20], and icosahedron projections[21, 22].
For depth estimation, equirectangular projection predominates. Omnidepth[13] introduced RectNet for efficient equirectangular feature extraction and depth prediction. ODE-CNN[14] proposed a hardware-software co-design system comprising: 1) a panoramic camera with binocular depth sensors, and 2) a depth estimation model leveraging RectNet baseline enhanced with spherical feature transformers between encoder-decoder stages and deformable convolutional spatial propagation for final prediction. Following vision transformers’ emergence, specialized models have proliferated: Panoformer[10] developed reference-point window self-attention for equirectangular features within an encoder-decoder architecture; Egformer[11], SGFormer[15], GLPanoDepth[23] subsequently incorporated global receptive fields and spherical geometry constraints. To address panoramic depth data scarcity, self-supervised methods [16, 17] operate directly on equirectangular RGB images.
@@ II-C Background Based Methods
Current state-of-the-art methods demonstrate strong quantitative performance across benchmarks. However, their 3D depth visualizations frequently exhibit structural inaccuracies in room geometry and over-smoothed corners. To address this limitation, recent approaches[32] integrate room structural priors into panoramic depth estimation. The core challenge lies in efficiently computing accurate background depth maps, which require precise room layout estimation.
Indoor panoramic layout estimation aims to detect room wall boundaries from input panoramas. Existing research predominantly adopts the Manhattan World assumption. LayoutNet[33] directly predicts per-pixel corner and boundary probability maps. Conversely, Dula-Net[34] decodes panoramas into equirectangular and perspective views, extracting wall-floor/ceiling boundaries via semantic segmentation. EquiConv[35] employs specialized convolutions to generate corner/edge probability maps. These early methods share a 2D segmentation paradigm, outputting full-resolution probability maps—an inefficient approach given layout estimation only requires boundary/corner identification. HorizonNet[36] thus reformulates the task as 1D sequence prediction using Bi-LSTMs, while HoHoNet[29] employs multi-head self-attention for similar 1D representation. LED-Net[37] further predicts layouts from horizontally compressed representations. Subsequently, AtlantaNet[38] projects panoramas onto ceiling/floor planes to avoid occlusion, predicting layouts via amalgamated projections. DMH-Net[39] extends this by mapping to cubemap faces before predicting boundaries and corners.
A critical secondary challenge involves fusing background depth with panoramic depth predictions. The recent BGDNet[32] incorporates room geometry as background depth when predicting the final depth map. However, it needs pretrained HorizonNet and SAM[40] to extract the background depth map from input panorama image. In this paper, we adopt a multi-task learning approach to integrate room layout and segmentation predictions needed for background depth calculation within a single framework. We also design a background depth calculation strategy based on background segmentation and a fusion strategy for coarse and background depth guided by background segmentation weights. Our method is detailed in the following sections.
Figure 2: The structure diagram of our proposed room geometry guided depth estimation framework. In terms of model structure, our framework includes a shared panorama encoder and three task-corresponding decoders. Based on the obtained layout map, coarse depth map, and background-segmentation map, the proposed framework decode fine-grained depth prediction.
@@ III Methodology
The room structural regularized depth estimation framework proposed in this paper is shown in Fig.2. From the view of neural network architecture, the framework includes a shared panorama encoder and three task-corresponding decoder modules. Moreover, our framework contain a room geometry-based background depth resolving strategy and a background-segmentation-guided fusion mechanism. The panorama encoder extract the multi-stage feature sets $F=\{f^{2^{i+1}\times}_{i}|i=1,2,3,4\}$ from the panorama image. The decoder modules corresponding to the subsequent three tasks use these features as input to estimate the corresponding results. The proposed room geometry-based background depth resolving strategy calculate the background depth based on panorama depth, room layout, background segmentation predictions. Then, background-segmentation-guided fusion mechanism regularize the depth decoder’s prediction based on background depth map and background segmentation predictions. In the following subsections we will introduce each module in detail.
@@ III-A Architecture
Figure 3: The structure diagram of layout feature aggregation module and pixel feature aggregation module.
@@
Panorama Encoder: In this paper, the framework we proposed uses the backbone proposed by PanoFormer as a common feature encoder. The backbone consists of the panorama transformer block and convolution 2D layer based downsample layer. The panorama transformer block contains a window self-attention mechanism designed for the panorama imaging process and a feed-forward layer designed based on depthwise separable convolution[41]. The entire backbone consists of an input projection layer and $4$ stage blocks. The input projection layer consists of a convolution layer whose kernel size is $3\times 3$ with stride of $2$, a 2D batchnorm layer and a ReLU activation function. The subsequent $4$ stage blocks all consist of a panorama transformer block and a double-downsampling layer composed of a 2D convolutional network. Each stage block will output the corresponding double-downsampled feature map result.
Layout Decoder: Considering that layout estimation is a non-pixel-level task while depth estimation and background segmentation are both pixel-level tasks, and some existing work [42] also mentioned that different tasks may have different requirements for features of different scales. To this end, we design a specific layout FA module for the layout task to further process the features extracted by the panorama encoder. The structure diagram of the layout feature aggregation module for the layout task is shown in Fig.3 (a) . This module use the height compress convolution with spatial stride $(2,1)$ to compress the feature maps $f^{4\times}_{1},f^{8\times}_{2},f^{16\times}_{3},f^{32\times}_{4}$. Then the compressed feature map is flattened to obtain a feature vectors set $\hat{f}_{1},\hat{f}_{2},\hat{f}_{3},\hat{f}_{4}$. We then cat these flattened features together and, following the Hohonet[25] model design experience, use a multi-head self-attention module to extract global feature information from the flattened feature vectors and output the final feature vector $\hat{f}_{cat}$. Based on the output features of the layout FA module, the layout decoder uses boundary and corner heads, consisting of a convolution 1D layer, a ReLU layer, and a batchnorm 1D layer, to predict the final room layout $S_{room}$.
Depth Decoder: Considering that both depth estimation and background segmentation tasks are pixel-level tasks, we follow the design idea of Hohonet and let them use a common pixel-level feature aggregation to process the multi-stage panorama feature. The structure of the pixel-level feature aggregation module we use is shown in Fig. 3 (b). This module consists of a panorama transformer block and an upsampling module. Given multi-scale feature map set $f^{1}_{4\times},f^{2}_{8\times},f^{3}_{16\times},f^{4}_{32\times}$, this module’s feature integration process can be described as:
$\hat{f}=\prod^{3}_{i=1}[UP(PT(f^{2^{i}\times}_{i}))+f^{2^{i+1}\times}_{i+1}]$ (1)
On the integrated multi-scale feature $\hat{f}$, the depth estimation decoder predict the coarse depth maps $S^{p}_{depth}=\{d_{ij}|i=1,2,...,W;j=1,2,...,H\}$.
background-Segmentation Decoder: As mentioned above, the background segmentation task and the depth estimation task share a pixel-level feature aggregation module. The structure diagram of the pixel-level feature aggregation module is shown in Fig. 3(b). Therefore, in our framework, the background-segmentation decoder itself has only one prediction head consisting of a convolution 2D layer, a ReLU function, and a batchnorm 2D layer. This prediction head predicts the corresponding 2D background segmentation result $S_{seg}=\{0\leq p_{ij}\leq 1|i=1,2,...,W;j=1,2,...,H\}$ based on the integrated multi-scale feature $\hat{f}$.
III-B Room Geometry-based Background Depth Resolving Strategy
@@
Given a room layout estimation prediction $S_{room}\in R^{W\times 3}$ to resolve the corresponding background depth, we need to obtain the camera-to-ground distance information. Existing methods often assume a fixed camera height when calculating background depth based on layout information. This assumption is stable and reliable in virtual rendering datasets, but it is not necessarily reliable when collecting data in the real world using a panoramic depth camera. To obtain more stable and reliable camera height information for each scene, we first designed a camera height resolution strategy.
As shown in the Fig. 4, define the coordinates of the camera origin $O$, and draw a normal line from the camera plane to the image plane, which intersects the wall plane at point $P$. Then draw a straight line perpendicular to the camera plane at point $P$, which intersects with the ceiling-wall boundary line in the room layout prediction at point $A$, and intersects with the floor-wall boundary line at point $B$. What our camera height calculation strategy needs to do is to calculate the camera height in the Fig. 4. Our method use the predictions of the depth decoder and room layout decoder as the input of camera height calculating method. The room layout predictions are responsible for providing the image plane coordinates of the intersection points $A$ and $B$ in the panorama image and depth maps. Based on the results of the depth maps prediction, we can get the distance between the corresponding line segments $OA$ and $OB$. On this basis, the distance from the camera to the ceiling and the camera height can be conveniently calculated using the following formula:
$|AP|=d_{c}\times sin(\phi_{c})$ (2)
$|PB|=d_{f}\times sin{\phi_{f}}$
The corresponding angles $\phi_{c}$ and $\phi_{f}$ can be calculated using the geometric imaging principle of a panoramic camera. Assuming that the vertical coordinate of point $A$ is $u_{ceil}$ and the vertical coordinate of point $B$ is $u_{floor}$, the calculation formulas for the two angles can be obtained using the following formulas:

### v2

# PAGCNet: A Pose-Aware and Geometry Constrained Framework for Panoramic Depth Estimation

@@ Abstract
Explicitly modeling room background depth as a geometric constraint has proven effective for panoramic depth estimation. However, reconstructing this background depth for regular enclosed regions in a complex indoor scene without external measurements remains an open challenge. To address this, we propose a pose-aware and geometry-constrained framework for panoramic depth estimation. Our framework first employs multiple task-specific decoders to jointly estimate room layout, camera pose, depth, and region segmentation from a input panoramic image. A pose-aware background depth resolving (PA-BDR) component uses tasks decoder’s prediction to resolve the camera pose. Subsequently, the proposed PA-BDR component uses the camera pose to compute the background depth of regular enclosed regions and uses this background depth as a strong geometric prior. Based on the output of the region segmentation decoder, a fusion mask generation (FMG) component produces a fusion weight map to guide where and to what extent the geometry-constrained background depth should correct the depth decoder’s prediction. Finally, an adaptive fusion component integrates this refined background depth with the initial depth prediction, guided by the fusion weight. Extensive experiments on Matterport3D, Structured3D, and Replica datasets demonstrate that our method achieves significantly superior performance compared to current open-source methods. Code is available at https://github.com/emiyaning/PAGCNet.
@@ I Introduction
To enable 3D understanding of indoor environments from a single omnidirectional image, depth estimation becomes a fundamental requirement[19]. This potential for inferring entire scene structure from a single panorama has motivated active research on panoramic depth estimation.
Figure 1: This figure presents 3D visualizations of the ground truth depth maps for two rooms with an atypical structure, shown from three distinct viewpoints: frontal, side, and bird’s-eye view (BEV). The first row depicts a room whose overall volume approximates a triangular prism shape. The second row illustrates a KTV room, in which the sofa cushions and the entire wall are merged into a single continuous structure.
Existing methods for monocular panoramic depth estimation primarily adopt three technical strategies to address inherent object distortions[10]. One line of work[25, 38] designs specialized feature extractors, informed by panoramic imaging principles, to directly process distorted equirectangular images. Another common strategy [32] projects the equirectangular panorama (ERP) onto another projection, extracts features separately from both, and then fuses them to predict the depth. A more recent direction[6] explicitly incorporates geometric priors; it uses estimated room layout to construct a background depth model, which significantly improves accuracy in geometrically discontinuous regions. While these methods achieve strong performance on benchmarks with regular, Manhattan-aligned layouts, they often overlook the prevalence of complex shaped rooms in real-world scenarios.
As illustrated in Fig. 1, real-world rooms are often characterized by irregular, non-Manhattan layouts. In contrast, the background depth estimated by BGDNet [6] based on its predicted room layout typically assumes a regular room structure and known camera pose, which limits the method’s performance in the scenarios depicted in Fig. 1. Nevertheless, prior work [31] has suggested that any irregular room can be decomposed into regular enclosed regions and irregular regions. The room layout annotations provided by various datasets usually correspond to the regular enclosed regions in such scenarios. Hence, the reliable identification and separation of regular and irregular regions in specifically structured indoor scenes, as well as the subsequent reconstruction of background depth for regular enclosed regions without external measurements, remains an open challenge.
To address existing problems in panoramic depth estimation, this paper proposes a depth estimation framework named as pose-aware and room geometry-constrained depth estimation framework (PAGCNet). The proposed framework based on multi-task learning unifying four tasks: depth estimation, camera pose estimation, room layout estimation, and region segmentation. Specifically, region segmentation decoder perform two semantic segmentation task to extract the irregular region mask and background mask from input panorama. The irregular region mask used as an indicator to tell depth estimator which region fall outside the enclosed background region defined by room layout, while background mask used to tell our framework to what extent the geometry-constrained background depth should correct the depth decoder’s prediction. Building on this output, our PA-BDR component optimize the camera pose decoder’s prediction and further resolving the enclosed region background depth. Then, our FGM component produces the fusion weight map according to the two mask prediction of region segmentation decoder. Finally, an adaptive fusion component integrates the enclosed region background depth with the depth decoder’s prediction under the guidance of fusion weight map, yielding a refined depth prediction. Extensive experimental results on the Matterport3D, Structured3D, Replica dataset show that our proposed methods can achieve significantly superior performance than current open-source methods. In general, our contributions can be summarized into the following three points:
  • We proposes PAGCNet, which calculates the enclosed region background depth to adaptively optimizes the final depth prediction.
  • We design a PA-BDR component to compute the background depth without external measurement of camera pose.
  • We introduce an FMG component and an adaptive fusion component. The FMG component uses the region segmentation decoder’s output to compute the fusion weight, while the adaptive fusion component uses this fusion weight to perform the fusion between background depth and depth decoder’s prediction.
@@ II-A Single Projection Inputs
Panoramic images employ spherical representations with $180^{\circ}$ vertical and $360^{\circ}$ horizontal fields of view. Processing typically involves projecting these spherical images onto 2D planes through perspective mapping. Common modalities include equirectangular[44, 7, 25, 38, 40], tangent[8, 18, 24], and icosahedron projections[39, 15].
For depth estimation, equirectangular projection predominates. Omnidepth[44] introduced RectNet for efficient equirectangular feature extraction and depth prediction. ODE-CNN[7] proposed a hardware-software co-design system. Following vision transformers’ emergence, specialized models have proliferated: Panoformer[25] developed reference-point window self-attention for equirectangular features within an encoder-decoder architecture; Egformer[38], SGFormer[40], GLPanoDepth[3] subsequently incorporated global receptive fields and spherical geometry constraints.
@@ II-C Background Based Methods
Current state-of-the-art methods demonstrate strong quantitative performance across benchmarks. However, their 3D depth visualizations frequently exhibit structural inaccuracies in room geometry and over-smoothed corners. To address this limitation, recent approaches[6] integrate room structural priors into panoramic depth estimation. Driven by advances in room layout estimation, current models[28, 42, 23, 33, 1, 31] demonstrate high accuracy in predicting layouts for rooms with regular, Manhattan-aligned structures. Yet, in real-world indoor images, room structures are often irregular and not based on a Manhattan layout, and the camera’s pose itself cannot be obtained in real time. This severely limits its performance in real-world indoor scenes with irregular structures. To eliminate the dependence on these two assumptions, our multi-task learning framework incorporates camera pose learning and irregular region segmentation tasks. Our method is detailed in the following sections.
Figure 2: The structure diagram of our proposed PAGCNet. In terms of model structure, our framework includes a shared panorama encoder and four task-corresponding decoders. Based on the obtained layout estimation, first stage depth, camera pose, irregular region mask and background region mask, the proposed framework decode fine-grained depth prediction.
@@ III Methodology
The proposed PAGCNet’s diagram in this paper is shown in Fig.2. From the view of network architecture, the framework includes a shared panorama encoder and four task-corresponding decoder modules. Moreover, our framework contain a PA-BDR component, a FGM component and an adaptive fusion component. The panorama encoder extract the multi-stage feature sets $F=\{f^{2^{i+1}\times}_{i}|i=1,2,3,4\}$ from the panorama image. The proposed PA-BDR component optimize the camera pose and resolving regular enclosed region’s background depth based on tasks decoder’s prediction. Then, FMG component extract the fusion weight map from input panorama with two sub binary semantic segmentation task’s prediction. Finally, the adaptive fusion component will fuse the regular enclosed region’s background depth and coarse depth prediction of depth decoder to get the refined depth predictions. In the following subsections we will introduce each module in detail.
@@ III-A Architecture
Panorama Encoder: In this paper, the framework we proposed uses the backbone proposed by PanoFormer[25] as a common feature encoder. The backbone consists of the panorama transformer block and convolution 2D layer based downsample layer. The panorama transformer block contains a window self-attention mechanism designed for the panorama imaging process and a feed-forward layer designed based on depth-wise separable convolution[12]. The entire backbone consists of an input projection layer and $4$ stage blocks. The input projection layer consists of a convolution layer whose kernel size is $3\times 3$ with stride of $2$, a 2D batch-norm layer and a ReLU activation function. The subsequent $4$ stage blocks all consist of a panorama transformer block and a double-downsampling layer composed of a 2D convolutional network. Each stage block will output the corresponding double-downsampled feature map result.
Layout Decoder: Considering that layout and camera pose estimation are non-pixel-level task while depth estimation and region segmentation are both pixel-level tasks. Some works[41] mentioned that different tasks may have different requirements for features of different scales. Following the hohonet[29], we first employ a height compression module to project multi-scale panoramic features from 2D into a 1D sequential representation. This sequence is then processed by the Transformer encoder introduced in Hohonet[29], and finally passed through a lightweight MLP to predict the enclosed room layout $S_{room}$.
Camera Pose Decoder: The camera pose decoder follows a feature processing pipeline consistent with that of the layout decoder: multi-scale 2D panoramic features are first compressed along the height dimension into a 1D sequence, which is then processed by a Transformer encoder. However, due to the distinct nature of the camera pose estimation task, the corresponding feature processing blocks in this decoder do not share weights with those in the layout decoder. After feature encoding, the camera pose is regressed through a lightweight multi-layer perceptron (MLP) head.
Depth Decoder: In this section, we maintain consistency with the original PanoFormer[25] paper. This part of the network consists of panorama transformer blocks and linear interpolation upsampling layers. The decoder network is structurally symmetrical to the panorama encoder. On the last panorama transformer block’s feature output $\hat{f}$, the depth estimation decoder predict the coarse depth maps $S^{p}_{depth}=\{d_{ij}|i=1,2,...,W;j=1,2,...,H\}$.
Region Segmentation Decoder: Given that the panorama encoder and depth decoder together constitute a U-Net-like architecture—commonly employed in semantic segmentation—we do not introduce separate feature extraction modules for the region segmentation decoder. Instead, we directly utilize the feature map $\hat{f}$ output by the final transformer block in the depth decoder. Two independent convolutional layers are then applied to this feature map to predict the irregular region mask $S_{ir}=\{p_{ij}\in{0,1}|i=1,2,...,W;j=1,2,...,H\}$ and the background segmentation mask $S_{seg}=\{0\leq p_{ij}\leq 1|i=1,2,...,W;j=1,2,...,H\}$, respectively.
III-B Pose Aware Background Depth Resolving Component
@@
III-B1 Camera Height Optimize
Given a room layout estimation prediction $S_{room}\in R^{W\times 3}$ to resolve the corresponding background depth, we need to obtain the camera-to-ground and camera-to-ceiling distance information. Existing methods often assume a fixed camera height or camera height can be obtained easily. This assumption is stable and reliable in virtual rendering datasets, but it is not necessarily reliable when collecting data in the real world using a panoramic depth camera. To obtain a stable and reliable camera height, our framework first use a camera pose decoder to predict an init camera height $\hat{h}_{c}$ from input panorama. Then, the PA-DBA component use the depth, room layout decoder’s prediction to calculate a camera height $\tilde{h}_{c}$. Finally, we count the mean value of $\hat{h}_{c}$ and $\tilde{h}_{c}$ to solve the accurate camera height $h_{c}$. In the following, we show the detail of calculation process of $\tilde{h}_{c}$.
From the results of room layout and region segmentation, we can obtain the foreground pixel coordinates of the upper and lower boundaries of the wall without occlusion. Here, we take a straightforward and easy-to-understand example (as the Fig. 3) to explain how to calculate our camera height $\tilde{h}_{c}$ from these occlusion free wall boundaries pixel coordinates.
As shown in Fig. 3, we define the origin of the camera plane as $O$, and then draw a horizon line on the camera plane centered at point $O$. Based on the results of the region segmentation decoder, we can select the portion of horizon line without occlusion and falling in the wall. We randomly select a point $P$ from these segment regions; the length of $OP$ is then the depth value of the corresponding pixel in the coarse depth map. Simultaneously, based on the horizontal coordinates of point $P$, we can obtain the two points $A$ and $B$ above and below the corresponding wall boundary line. The lengths of $OA$ and $OB$ can also be obtained from the corresponding depth map pixels $d_{c}$ and $d_{f}$. Using basic geometry, we can calculate the camera height $AP$ and the height $BP$ from the camera to the ceiling. The calculation formulas are as follows:
$|AP|=d_{c}\times sin(\phi_{c}),$ $\ |PB|=d_{f}\times sin{\phi_{f}},$ (1)
where angles $\phi_{c}$ and $\phi_{f}$ can be calculated using the geometric imaging principle of a panoramic camera. Assuming that the vertical coordinate of point $A$ is $u_{ceil}$ and the vertical coordinate of point B is $u_{floor}$, the calculation formulas for the two angles can be obtained using the following formulas: